The resultant of the three concurrent forces acting on the eyebolt is the force R = 800j lb. Determine the magnitude of the force P and the angle θ that specifies the direction of the 900-lb force.


1
Expert's answer
2021-06-07T09:40:47-0400

Assume the body diagram below



For equilibrium condition in y-direction

"\Sigma F_y=0"

"900sin\theta+500sin\alpha=800"

"900sin\theta+500(3/5)=800"

"sin\theta=500/900=5/9"

"\theta=33.74^o"

for the horizontal force, x-direction

"\Sigma F_x=0"

"-900cos\theta+500cos\alpha+P=0"

"-P=-900cos(33.74^o)+500(4\/5)"

"P=348.40N"